Consider the following information about an economy with a labor force

participation rate of 80%.

Labor force 500

Number of people employed 425

Number of discouraged workers 15

a. What is the population 15 years and older of this economy?

b. How many people are unemployed? What is the unemployment rate? .

c. If the discouraged workers joined the labor force, but total employment did

not change, what would be the new unemployment rate?


1
Expert's answer
2021-07-05T08:27:29-0400

a. What is the population 15 years and older of this economy?

LFPR = "\frac{labor force}{civilian noninstitutional population} \times100"


80% = "\frac{500}{x} \times100"

0.8x = 500

x = 625

The population 15 years and older is 625


b. How many people are unemployed? What is the unemployment rate? .

The number of unemployed = total labor force - employed people

= 500 - 425

= 75


Unemployment rate = "\frac{number of unemployed}{total labor force} \times100"

= "\frac{75}{500} \times100"


unemployment rate = 15%


c. If the discouraged workers joined the labor force, but total employment did

not change, what would be the new unemployment rate?

The new unemployment rate would be:

"\frac{(75+15)}{(500+15)} \times100"


"\frac{90}{515} \times100"

= 17.5%